# Inference-time Scaling of Diffusion Models through Classical Search

### v1

@@ Abstract
Classical search algorithms have long underpinned modern artificial intelligence. In this work, we tackle the challenge of inference-time control in diffusion models—adapting generated outputs to meet diverse test-time objectives—using principles from classical search. We propose a general framework that orchestrates local and global search to efficiently navigate the generative space. It employs a theoretically grounded local search via annealed Langevin MCMC and performs compute-efficient global exploration using breadth-first and depth-first tree search. We evaluate our approach on a range of challenging domains, including planning, offline reinforcement learning, and image generation. Across all tasks, we observe significant gains in both performance and efficiency. These results show that classical search provides a principled and practical foundation for inference-time scaling in diffusion models. Project page at diffusion-inference-scaling.github.io.
@@ 1 Introduction
Classical search algorithms have laid the foundation for modern artificial intelligence 10.5555/1671238 . In discrete settings, graph search algorithms are widely used to explore the state space. Breadth-first search (BFS) moore1959shortest and depth-first search (DFS) tarjan1972depth traverse the search tree in a fixed order. To better leverage problem-specific information, best-first search methods pearl1984heuristics , such as A* hart1968formal , use a heuristic to evaluate and prioritize states. Alternatively, local search methods, such as hill-climbing (10.5555/1671238, , Sec. 4.1), explore neighboring states. More recent techniques like gradient descent and Markov Chain Monte Carlo (MCMC) have become widely adopted in optimization and probabilistic inference, underpinning many modern AI models.
@@
Diffusion models ho2020denoising have shown impressive performance in generative modeling for continuous domains such as images dhariwal2021diffusion , videos ho2022video , and world modeling yanglearning . They are also increasingly used in robotics and decision-making liu2024rdt ; black2024pi_0 ; team2024octo to generate diverse actions chi2023diffusion . However, generated samples may not always align with physical laws song2023loss or human intent wallace2024diffusion , and the vast generative space often necessitates multiple trials to produce satisfactory outputs xie2025sana . To address this, we scale up inference-time compute using strategic search methods that navigate the generative manifold for high-quality samples. We formalize sample evaluation using a verifier function $f({\bm{x}}_{0})$ defined on ground truth samples, which measures the quality of the sample. Such verifiers could be reward functions xu2023imagereward , Q-functions lu2023contrastive , classifier conditions $p(c|{\bm{x}}_{0})$ ye2024tfg ; dhariwal2021diffusion , and multi-modal LLMs huang2023t2icompbench .
@@
To efficiently search the generative space induced by diffusion, we revisit classical search principles. Inspired by local search, when the verifier $f$ is differentiable, we use Langevin MCMC to explore neighborhoods of current samples, guided by the verifier gradient and “score function” vincent2011score_matching from the diffusion model. To explore diverse modes in the complex distribution generated by diffusion models, we search the Markov chain of the “denoising process” with global search methods. Viewing sampling as a search tree, we use BFS and DFS to gradually explore the states while denoising. Like best-first search methods, we can evaluate intermediate states ${\bm{x}}_{t}$ using the verifier $f({\bm{x}}_{0|t})$, expanding high-quality paths more effectively. An illustration of local and global search is shown in Fig. 1.
Recently, Ma et al. ma2025inference also explored inference-time scaling of diffusion models using search-based techniques, focusing on image generation tasks. Their zero-order search approach can be interpreted as a instance of local search, while their path-based strategies a combination of both local and global search. However, their method does not effectively leverage verifier gradients, and we introduce a local search strategy with annealed Langevin MCMC, efficiently steering sampling toward high-score regions. Moreover, we systematically evaluate the compute efficiency of search methods. Finally, we extend inference-time scaling beyond image generation to a broader set of domains.
To this end, we summarize our key contributions as follows:
@@
    We propose a general and unified framework for efficient inference-time search in diffusion models, drawing inspiration from classical search algorithms. Specifically, we introduce a local search method based on Langevin MCMC that theoretically unifies prior gradient-based approaches, and we develop efficient global search strategies using tree-based algorithms such as BFS and DFS.
@@
    We implement inference-time scaling across a range of tasks—including planning, offline reinforcement learning, and image generation—and introduce task-specific verifiers. Our results demonstrate that inference-time scaling with search against verifiers can significantly improve the performance on many downstream tasks. Throughout the experiments, we show that our proposed approaches consistently surpass baselines in terms of compute efficiency.
Figure 1: Illustration of our search framework. Bottom left: direct sampling results in samples with low verifier scores. Bottom middle: Local search navigates samples to nearby high-score regions, but suffers from local maximums and OOD samples. Bottom right: introducing global search explores diverse modes globally, allocating more compute to high-quality samples, and sampling from global optimal modes.
@@ 2 Related Works
Here we provide a brief overview of inference-time scaling and diffusion models. For a more in-depth literature review see Appendix. B.
Scaling inference-time compute has its roots in cognitive science, often associated with “System 2” thinking kahneman2011thinking ; anthony_thinking_2017 . Yao et al. yao2023tree conceptualize problem-solving in large language models (LLMs) as a combinatorial tree search problem newell1959report ; newell1972human , proposing the Tree-of-Thoughts framework, which leverages breadth-first and depth-first search for more efficient reasoning.
Inference-time scaling has recently shown impressive performance with LLMs in reasoning and problem-solving tasks. For instance, jaech2024openai ; van2016deep train models to generate long chain-of-thought (CoT) wei2022chain , while others such as snell2024scaling ; wu2024inference incorporate process reward models lightman2023let to guide the generation process more effectively. These works focus on efficient exploration in discrete search spaces, often employing global search strategies. In contrast, diffusion models operate in a continuous, multimodal generative space. Furthermore, the autoregressive nature of LLMs allows for efficient evaluation of intermediate generation steps—something not directly feasible in diffusion models dhariwal2021diffusion ; lu2023contrastive .
Inference-time scaling for diffusion models remains relatively underexplored. Recent works such as ma2025inference ; nakamotosteering introduce BFS-style global search algorithms, resampling candidate samples (or “particles”) at each noise level based on verifier scores (potentials). However, these methods do not effectively perform local search. Alternatively, Du et al. du_learning_2024 propose learning a series of annealed energy landscapes and applying Langevin MCMC for iterative refinement. In this work we propose a general principled framework, encompassing prior methods as instances of local search and global search.
@@ 3.1 Diffusion Probabilistic Models
Suppose we have $D$-dimensional random variable ${\bm{x}}_{0}\in\mathbb{R}^{D}$ with distribution $p_{0}({\bm{x}}_{0})$. Diffusion models ho2020denoising ; song2020denoising and the more general flow models lipman2022flow ; albergo2022building are generative models that turn noise into data via a stochastic process $\left\{{\bm{x}}_{t}\right\}_{t=0}^{T}$. The forward “noising” process with $t>s$ can be defined as:
@@
where $\alpha_{t}$, $\sigma_{t}$ are referred as the noise schedule with $\alpha_{0}=\sigma_{T}=1$, $\alpha_{T}=\sigma_{0}=0$. We can thus write the random variables ${\bm{x}}_{t}$ as an interpolation between data and noise ma2024sit :
@@
and denote $q_{t}({\bm{x}}_{t})$ as the marginal distribution of ${\bm{x}}_{t}$. To model the reverse “denoising” process, we train the model using the denoising objective ho2020denoising :
@@
which is equivalent of learning the score function of $q_{t}({\bm{x}}_{t})$ vincent2011score_matching , as the ground truth of $\epsilon_{\theta}({\bm{x}}_{t},t)$ is $-\sigma_{t}\nabla_{{\bm{x}}_{t}}\log q_{t}({\bm{x}}_{t})$. To generate samples, we transform noise into data via the reverse transition kernel $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$. In practice, we either sample ${\bm{x}}_{t-1}$ using deterministic samplers like DDIM song2020denoising :
@@
or stochastic samplers like DDPM ho2020denoising ; nichol2021improved :
@@ 3.2 Compositional and Controllable Generation of DPMs
Given a base diffusion model with data distribution $p_{0}({\bm{x}}_{0})$, one may wish to sample ${\bm{x}}_{0}$ with some constraints or conditions $f({\bm{x}}_{0})$. Exact diffusion sampling from the composed distribution $\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})$ would require training a time-dependent $f$ on data generated by $p_{0}$ dhariwal2021diffusion ; lu2023contrastive , which may not be applicable in practice. Thus, we adopt optimization based methods to approximate the target distribution.
@@
When sampling from a compositional distribution composed of multiple probability distributions, $\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})\hat{p}_{0}({\bm{x}}_{0})$, du2023reduce proposes annealed Langevin MCMC sampling. In this approach, a sequence of annealed distributions $\tilde{q}_{t}({\bm{x}}_{t})\propto q_{t}({\bm{x}}_{t})\hat{q}_{t}({\bm{x}}_{0})$ is constructed, and samples are drawn using Langevin dynamics welling2011bayesian :
@@
where ${\bm{x}}_{0|t}=\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]=\frac{{\bm{x}}_{t}-\sigma_{t}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)}{\alpha_{t}}$. This method approximates the intractable posterior with the posterior mean: $\mathbb{E}_{{\bm{x}}_{0}|{\bm{x}}_{t}}[f({\bm{x}}_{0})]\approx f(\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}])$. To enhance the guidance strength, yu2023freedom propose to apply a recurrence strategy, which first samples ${\bm{x}}_{t-1}$ via $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$, add the guidance gradient, then add noise back to ${\bm{x}}_{t}$ through the forward process $q_{t}({\bm{x}}_{t}|{\bm{x}}_{t-1})$:
@@ 4 Methods
Given a pretrained diffusion model $\epsilon_{\theta}({\bm{x}}_{t},t)$ with a base distribution $p_{0}({\bm{x}}_{0})$, at test-time, we often wish to optimize the generation process to satisfy task-specific objectives. For instance, in reinforcement learning, this might involve generating high-value actions; in image synthesis, we may seek images that satisfy certain constraints such as color or shape; and in trajectory generation, we may want the outputs to conform to physical laws. In this paper, we are interested in how to scale test-time inference to follow such objectives.
@@
$\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})^{\lambda}\,,$
@@
Since exact sampling from the distribution is often impractical, we aim to search the manifold for the target samples at inference time, both locally and globally. First, we search the vicinity of the sample using hill-climbing style local search methods, guided by the verifier gradient. However, local search may stuck in a local maximum, where we then propose global graph search algorithms to explore the diverse modes in the complex generative landscape of diffusion models.

### v2

@@ Abstract
Classical search algorithms have long underpinned modern artificial intelligence. In this work, we tackle the challenge of inference-time control in diffusion models—adapting generated outputs to meet diverse test-time objectives—using principles from classical search. We propose a general framework that orchestrates local and global search to efficiently navigate the generative space. It performs compute-efficient global exploration using breadth-first and depth-first tree search and employs a theoretically grounded, scalable local search via annealed Langevin MCMC. We evaluate our approach on a range of challenging domains, including planning, offline reinforcement learning, and image generation, and observe significant gains in both performance and efficiency over baseline methods. These results demonstrate that classical search offers a principled and practical foundation for inference-time scaling in diffusion models. By jointly scaling local and global search for the first time, our framework establishes a new Pareto frontier across challenging decision-making domains. Project page at diffusion-inference-scaling.github.io.
@@ 1 Introduction
Classical search algorithms have laid the foundation for modern artificial intelligence (Russell & Norvig, 2009). In discrete settings, graph search algorithms are widely used to explore the state space. Breadth-first search (BFS) (Moore, 1959) and depth-first search (DFS) (Tarjan, 1972) traverse the search tree in a fixed order. To better leverage problem-specific information, best-first search methods (Pearl, 1984), such as A* (Hart et al., 1968), use a heuristic to evaluate and prioritize states. Alternatively, local search methods, such as hill-climbing (Russell & Norvig, 2009, Sec. 4.1), explore neighboring states. More recent techniques like gradient descent and Markov Chain Monte Carlo (MCMC) have become widely adopted in optimization and probabilistic inference, underpinning many modern AI models.
@@
Diffusion models (Ho et al., 2020) have shown impressive performance in generative modeling for continuous domains such as images (Dhariwal & Nichol, 2021), videos (Ho et al., 2022). They are also increasingly used in robotics and decision-making (Liu et al., 2024; Black et al., 2024; Team et al., 2024) to generate diverse actions (Chi et al., 2023). However, generated samples may not always align with physical laws (Song et al., 2023) or human intent (Wallace et al., 2024), and the vast generative space often necessitates multiple trials to produce satisfactory outputs (Xie et al., 2025). To address this, we scale up inference-time compute using strategic search methods that navigate the generative manifold for high-quality samples. We formalize sample evaluation using a verifier function $f({\bm{x}}_{0})$ defined on ground truth samples, which measures the quality of the sample. Such verifiers could be reward functions (Xu et al., 2023a), Q-functions (Lu et al., 2023), classifier conditions $p(c|{\bm{x}}_{0})$ (Ye et al., 2024; Dhariwal & Nichol, 2021), and VLMs (Huang et al., 2023).
@@
To efficiently search the generative space of diffusion models, we revisit classical search principles. To capture diverse modes in the complex distributions generated by diffusion models, we view sampling as traversing a search tree, employing BFS and DFS to progressively explore states during denoising. Similar to best-first search, we evaluate intermediate states ${\bm{x}}_{t}$ with the verifier $f({\bm{x}}_{0|t})$, prioritizing high-quality paths and efficiently allocating compute via branching and backtracking. To go beyond the base model and obtain higher-quality samples, we perform local search via Langevin MCMC, exploring the neighborhood of current samples under guidance from both the verifier gradient and the diffusion model’s “score function” (Vincent, 2011). By jointly optimizing the compositional objective of the diffusion model and the verifier (Du et al., 2023), our local search surpasses the capabilities of the base model. An overview of our framework is shown in Fig. 1.
Recent works scale diffusion model inference via particle-based SMC (Kim et al., 2025; Singhal et al., 2025a; Wu et al., 2023) and tree-based methods (Guo et al., 2025), typically as BFS with fixed schedules. We generalize these with a BFS-based framework, clarifying prior design choices and providing an improved BFS surpassing prior methods. Inspired by DFS, we add adaptive backtracking to allocate compute adaptively, surpassing BFS in efficiency and adaptivity. While global search remains limited to base-distribution modes, scaling local search with Langevin MCMC explores high-reward regions beyond the model, proving effective in challenging decision-making tasks. Finally, we jointly scale local and global search with the classical AI search principles, demonstrating superior performance over prior methods, which scale local and global search in isolation.
Our key contributions are summarized as follows:
@@
    For global tree search, we elucidate the design space of prior BFS-style methods and provide an improved BFS design. We further present the first adaptive DFS algorithm for diffusion inference scaling, offering superior efficiency and adaptivity.
@@
    We introduce a theoretically grounded local search method using annealed Langevin MCMC, demonstrating superior performance in challenging decision-making domains.
  • iii)
    We propose a unified framework for inference scaling that integrates both global and local search within the classical AI search paradigm. By jointly scaling local and global search for the first time, we advance the Pareto frontier of inference-time scaling for diffusion models across diverse domains.
Figure 1: Illustration of our search framework. Bottom left: direct sampling results in samples with low verifier scores. Bottom middle: global search identifies high score modes within the base distribution. Bottom right: local search further optimizes the samples for higher quality, driven by the gradient signal.
@@ 2 Related Works
Here, we provide a brief overview of inference-time scaling with diffusion models. For a more comprehensive literature review and discussion of concurrent works, see Appendix B.
Particle-based scaling methods
Recent works (Kim et al., 2025; Singhal et al., 2025a) propose SMC-style particle filtering methods, scaling inference compute by increasing the number of particles and improving efficiency via resampling. Similarly, tree-search-based methods (Li et al., 2024; Guo et al., 2025; Ma et al., 2025) evaluate intermediate nodes and expand promising candidates. Both approaches can be seen as special cases of our BFS tree search framework, which denoise a set of particles in parallel, applying branching and filtering based on intermediate reward signals. Additional to linear BFS-style algorithms, we also propose DFS-style algorithms with non-linear adaptive backtracking, demonstrating superior performance over prior methods.
Gradient-based guidance methods
To sample from a conditional distribution, classifier guidance (Dhariwal & Nichol, 2021) utilizes the gradient from a trained noise-dependent classifier to compute the conditional score function (Song et al., 2020b). However, such noise-dependent classifier requires additional training and data collection. To use classifiers defined on clean samples, training-free guidance methods (Ye et al., 2024; Chung & Ye, 2022; Song et al., 2023; Yu et al., 2023; He et al., 2023) approximates noisy conditional probability using the denoised output ${\bm{x}}_{0|t}$. Such methods are inheritability biased due to their first order approximation. Different from prior methods that rely on the conditional diffusion process, we sample from the compositional distribution via annealed Langevin MCMC, which provides asymptotically exact sampling without any training.
@@ 3.1 Diffusion Probabilistic Models
Suppose we have $D$-dimensional random variable ${\bm{x}}_{0}\in\mathbb{R}^{D}$ with distribution $p_{0}({\bm{x}}_{0})$. Diffusion models (Ho et al., 2020; Song et al., 2020a) and the more general flow models (Lipman et al., 2022; Albergo & Vanden-Eijnden, 2022) are generative models that turn noise into data via a stochastic process $\left\{{\bm{x}}_{t}\right\}_{t=0}^{T}$. The forward “noising” process with $t>s$ can be defined as:
@@
where $\alpha_{t}$, $\sigma_{t}$ are referred as the noise schedule with $\alpha_{0}=\sigma_{T}=1$, $\alpha_{T}=\sigma_{0}=0$. We can thus write the random variables ${\bm{x}}_{t}$ as an interpolation between data and noise (Ma et al., 2024):
@@
and denote $q_{t}({\bm{x}}_{t})$ as the marginal distribution of ${\bm{x}}_{t}$. To model the reverse “denoising” process, we train the model using the denoising objective (Ho et al., 2020):
@@
which is equivalent of learning the score function of $q_{t}({\bm{x}}_{t})$ (Vincent, 2011), as the ground truth of $\epsilon_{\theta}({\bm{x}}_{t},t)$ is $\sigma_{t}\nabla_{{\bm{x}}_{t}}\log q_{t}({\bm{x}}_{t})$. To generate samples, we transform noise into data via the reverse transition kernel $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$. In practice, we either sample ${\bm{x}}_{t-1}$ using deterministic samplers like DDIM (Song et al., 2020a):
@@
or stochastic samplers like DDPM (Ho et al., 2020; Nichol & Dhariwal, 2021):
@@ 3.2 Compositional and Controllable Generation of DPMs
Given a base diffusion model with data distribution $p_{0}({\bm{x}}_{0})$, one may wish to sample ${\bm{x}}_{0}$ with some constraints or conditions $f({\bm{x}}_{0})$. Exact diffusion sampling from the composed distribution $\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})$ would require training a time-dependent $f$ on data generated by $p_{0}$ (Dhariwal & Nichol, 2021; Lu et al., 2023), which may not be applicable in practice. Thus, we adopt optimization based methods to approximate the target distribution.
@@
When sampling from a compositional distribution composed of multiple probability distributions, $\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})\hat{p}_{0}({\bm{x}}_{0})$, Du et al. (2023) proposes annealed Langevin MCMC sampling. In this approach, a sequence of annealed distributions $\tilde{q}_{t}({\bm{x}}_{t})\propto q_{t}({\bm{x}}_{t})\hat{q}_{t}({\bm{x}}_{0})$ is constructed, and samples are drawn using Langevin dynamics (Welling & Teh, 2011):
@@
where ${\bm{x}}_{0|t}=\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]=\frac{{\bm{x}}_{t}-\sigma_{t}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)}{\alpha_{t}}$. This method approximates the intractable posterior with the posterior mean: $\mathbb{E}_{{\bm{x}}_{0}|{\bm{x}}_{t}}[f({\bm{x}}_{0})]\approx f(\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}])$. To enhance the guidance strength, Yu et al. (2023) propose to apply a recurrence strategy, which first samples ${\bm{x}}_{t-1}$ via $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$, add the guidance gradient, then add noise back to ${\bm{x}}_{t}$ through the forward process $q_{t}({\bm{x}}_{t}|{\bm{x}}_{t-1})$:
@@ 4 Methods
Given a pretrained diffusion model $\epsilon_{\theta}({\bm{x}}_{t},t)$ with a base distribution $p_{0}({\bm{x}}_{0})$, at test-time, we often wish to optimize the generation process to satisfy task-specific objectives. For example, RL may require generating high-value actions, image synthesis may seek constraint-satisfying images, and trajectory generation may demand physically valid outputs. In this paper, we are interested in how to scale test-time inference to follow such objectives.
@@
$\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})^{\lambda}\,$, (5)
@@
Since exact sampling from the distribution is often impractical, we aim to search the manifold for the target samples at inference time, both globally and locally. First, we explore the diverse modes in the complex generative landscape of diffusion models through global graph search algorithms. However, global search alone can not generate samples beyond the pretrained model. We then propose to search the vicinity of the sample using hill-climbing style local search methods, guided by the verifier gradient. By combining local search with global search, we can avoid being stuck in local optima, advancing the performance of the base model by sampling beyond the pretrained modes.
4.1 Global Search for Mode Identification
To efficiently explore the modes of a diffusion model, we represent the Markov chain of its denoising process as a fixed-depth tree, where the transition kernel $\tilde{p}_{\theta}({\bm{x}}{t-1}|{\bm{x}}_{t})$ may correspond to either deterministic or stochastic samplers. This abstraction enables the use of classical tree-search heuristics to design compute-efficient exploration methods.
A simple baseline is best-of-$N$ (BoN) sampling: generate $N$ independent trajectories and select the one with the highest verifier score at the final step. While straightforward, this approach discards valuable information from intermediate stages. To improve efficiency, inspired by best-first search (Pearl, 1984), we evaluate intermediate particles using the denoised estimate ${\bm{x}}_{0|t}$. This allows dynamic allocation of computation based on particle quality, using techniques such as branching and backtracking. We next describe two representative efficient search strategies—breadth-first search (BFS) and depth-first search (DFS)—adapted for diffusion model sampling. By prioritizing the expansion of nodes with higher score estimates and backtracking from low-quality regions of the tree, we can more effectively navigate the generative space and identify high-quality modes. An illustration of the tree search methods is provided in Fig. 2.
4.1.1 Unified BFS-style linear search
Figure 2: Illustration of global tree search algorithms.
Inspired by breadth-first search (BFS), which expands nodes level by level, we denoise a set of particles in parallel at each noise level. We score each intermediate particle $\left\{{\bm{x}}_{t}^{k}\right\}_{k=1}^{N}$ using estimates of its verifier score $f({\bm{x}}_{0|t}^{k})$, and dynamically reallocate computation by sampling more children $n_{t}^{k}$ for high-scoring nodes. We provide a general design space for schedules, scoring functions, and resampling strategies that unifies previous tree-search-based and particle-based baselines such as SVDD (Li et al., 2024), DAS (Kim et al., 2025), and FK-steering (Singhal et al., 2025b). The pseudocode is shown in Alg. 3, and more details can be found in Sec. D.1.
Tempering
To mitigate score estimation bias in early steps, Kim et al. (2025) increases weights on smaller time steps with $\tau_{T}<\tau_{T-1}\cdots<\tau_{0}$, re-weighting scores with $\tau_{t}f({\bm{x}}_{0|t}^{k})$. Li et al. (2024) samples only from the top-scoring particle, i.e., $\tau_{t}=\infty$. We consider three different tempering design choices: $\textbf{Constant}:\tau_{t}=\tau,~\textbf{Increase}:\tau_{t}=\left((1+\gamma)^{T-t}-1\right)\tau,~\textbf{Inf}:\tau_{t}=\infty$~.
Scoring
Following Kim et al. (2025); Singhal et al. (2025b), we propose to evaluate intermediate particles $\widehat{f}({\bm{x}}_{t}^{k})$ using functions of the estimated rewards $f({\bm{x}}_{0|s}^{k})$ along the sampling path. This approach accounts for reward variation during sampling by considering not only the current reward but also its trajectory. Specifically, we define the following scoring functions $\widehat{f}({\bm{x}}_{t}^{k})$ based on $f({\bm{x}}_{0|s}^{k})$ for $s\geq t$: $\textbf{Current}:\tau_{t}f({\bm{x}}_{0|t}^{k}),~\textbf{Difference}:\tau_{t}f({\bm{x}}_{0|t}^{k})-\tau_{t+1}f({\bm{x}}_{0|t+1}^{k}),~\textbf{Max}:\max_{s\geq t}\tau_{s}f({\bm{x}}_{0|s}^{k}).$
Resampling
Given $\widehat{f}({\bm{x}}_{t}^{k})$, we allocate particles as $\left(n_{t}^{1},\cdots,n_{t}^{N}\right)=\text{Resample}\left(N;w_{t}^{1},\cdots,w_{t}^{N}\right)$, where $n_{t}^{k}$ is the number of children for ${\bm{x}}_{t}^{k}$ and $w_{t}^{k}=\text{softmax}{\widehat{f}({\bm{x}}_{t}^{k})}$ is the softmax score of particle $k$. Multinomial resampling (Singhal et al., 2025b) samples $n_{t}^{k}$ independently from the multinomial distribution $w_{t}^{k}$, but suffers from large variance. Kim et al. (2025) adopts the Srinivasan sampling process (SSP) resampling (Alg. 4) for reduced variance. We compare the baseline Multinomial resampling and the variance-reduced SSP (Kim et al., 2025); see Gerber et al. (2020) for detailed discussions of other resampling methods.
Prior methods are special cases of BFS: SVDD (Li et al., 2024) = BFS (Inf, Current, Multinomial); DAS (Kim et al., 2025) = BFS (Increase, Difference, SSP); FK-steering (Singhal et al., 2025a) = BFS (Constant, Max, Multinomial). Ablations (Sec. 5.1) show SSP resampling is key for performance, and our improved BFS (Increase, Max, SSP) consistently outperforms prior methods in efficiency.
4.1.2 DFS-style non-linear search
Depth-first search (DFS) explores one branch of the search tree as deeply as possible before backtracking. In our setting, this corresponds to iteratively denoising a single particle until its verifier score drops below a predefined threshold: $f({\bm{x}}_{0|t})\leq\delta_{t}$, where $\delta_{t}$ is a threshold representing the quality requirement of the users.
Once the constraint is violated, the algorithm backtracks by reintroducing noise, moving to a higher noise level $t_{\text{next}}=t+\Delta_{T}$ using the forward diffusion process $q({\bm{x}}_{t_{\text{next}}}|{\bm{x}}_{t})$ in Eq. 1. This allows the model to restart the denoising process from a different region of the manifold, encouraging exploration of diverse modes. Unlike the small noise injection and fixed schedule used in SoP (Ma et al., 2025) for local exploration, DFS performs global exploration with large backtracking depth $\Delta_{T}\geq\frac{T}{4}$, and adopts an score-adaptive exploration strategy .
A key strength of DFS is its ability to allocate compute adaptively: difficult prompts and low-quality trajectories naturally trigger more backtracking and exploration, while easier instances are solved more directly. This dynamic behavior is driven purely by the verifier signal, without needing to know the difficulty in advance as in Snell et al. (2024). Also, the threshold acts as a control knob for users to balance output quality and computation resources, where higher threshold automatically scales compute for better output. As shown in Sec. 5.2, this adaptive strategy leads to substantial gains in efficiency and performance over prior methods, and even our improved BFS implementation. The pseudocode is in Alg. 5 and more details can be found in Sec. D.2.